OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x8 + 3x6) cubic feet. What polynomial represents the height of the box in feet?

OpenStudy (campbell_st):

well the base area is \[A= x^2 \] multiply length by width and \[V = 4x^8+ 3x^6\] take a factor of x^2 out ot the volume \[V =x^2(4x^6 + 3x^4)\] now compare it to the volume of a prism V = base area x height you know the base area is x^2 so you should be able to identify the height of the prism

OpenStudy (anonymous):

is it = 4x + 3

OpenStudy (campbell_st):

V = base area x height you know the base area is x^2 \[V = x^2 \times (4x^6 + 3x^4)\] the height is the other factor... shown above

OpenStudy (anonymous):

4x6 + 3x4 yes , thanks